# A Hypernetwork-Based Approach to KAN Representation of Audio Signals

### v2

@@ 4.2. Hypernetworks
When using hypernetworks, we only consider NeRF, SIREN, and KAN as the target network architectures. We compare our method with the state-of-the-art model HyperSound (Szatkowski et al., 2023) and follow the same experimental setting. We use the VCTK dataset resampled to $22,050$ Hz. The inputs are truncated to length of 32,768 samples. The last 10 speakers are used as the validation set. All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch. We use AdamW optimizer (Loshchilov and Hutter, 2019) and OneCycleLR scheduler (Smith and Topin, 2018). The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is $1\mathrm{e}{-6}$.
@@ Results
[…]
@@
[…]
@@
[…]
@@ 4.3. KAN Ablation Study
[…]
@@
[…]
@@
[…]
@@ G Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
@@ 4.4. Frequency-based analysis
Spectrograms
Figure 3. The comparison of spectrograms. We compare the original sample (GT) with the waveforms reconstructed using different INR approaches.
@@
To demonstrate KAN’s improvements over previous INRs in high-frequency information of the reconstruction, we designed an additional experiment. We can treat the absolute values of the FFT of the audio signal as probability distribution of the base frequencies. Then, we can compare distributions of the ground truth and the reconstruction using Wasserstein distance. By using $[0,1,4,...(len(audio)-1)^{2}]$ as the support of the sample instead of intuitive $[0,1,2,...len(audio)-1]$, we can make the metric more biased towards high frequencies. Using this metric in our standard experimental setting for INRs, we obtain the following results, presented in Table 11. They indicate that KAN achieves the lowest score of 79.72 among all baseline methods. The second-best outcome is obtained by WIRE, while RFF performs the worst.
@@ A.1. Extended ablation study
[…]
@@
[…]
@@
In comparison to the 1.5 s case, where a spline order of size 3 is clearly the optimal hyperparameter setting, its advantage is less evident in Table 14. KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI metrics.
@@
[…]
@@
The results of the grid size parameter (displayed in Table 15) provide similar conclusions and tendencies as in the 1.5 s scenario (Table 10) — increasing the grid size leads to the improvement in all metrics but LSD.
@@
[…]
@@
[…]
@@
[…]

### v3

@@ 4.2. Hypernetworks
When using hypernetworks, we only consider NeRF, SIREN, and KAN as the target network architectures. We compare our method with the state-of-the-art model HyperSound (Szatkowski et al., 2023) and follow the same experimental setting. We use the VCTK dataset resampled to $22,050$ Hz. The inputs are truncated to length of 32,768 samples. The last 10 speakers are used as the validation set. All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch. We use AdamW optimizer (Loshchilov and Hutter, 2019) and OneCycleLR scheduler (Smith and Topin, 2018b). The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is $1\mathrm{e}{-6}$.
@@ Results
Network $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
Net. $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@ 4.3. KAN Ablation Study
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@ G Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@ 4.4. Frequency-based analysis
To demonstrate KAN’s improvements over previous INRs in high-frequency information of the reconstruction, we designed an additional experiment. We can treat the absolute values of the FFT of the audio signal as probability distribution of the base frequencies. Then, we can compare distributions of the ground truth and the reconstruction using Wasserstein distance. By using $0,1,4,...(len(audio)-1)^{2}]$ as the support of the sample instead of intuitive [0,1,2,len(audio)1][0,1,2,...len(audio)-1], we can make the metric more biased towards high frequencies. Using this metric in our standard experimental setting for INRs, we obtain the following results, presented in Table 11. They indicate that KAN achieves the lowest score of 79.72 among all baseline methods. The second-best outcome is obtained by WIRE, while RFF performs the worst.
@@
Figure 3. The comparison of spectrograms. We compare the original sample (GT) with the waveforms reconstructed using different INR approaches.:
Spectrograms
@@ 5. Conclusion
Acknowledgements.
This work was funded by several institutions. P. Marszałek was supported by the National Center of Science (Poland) Grant No.
2023/50/E/ST6/00169. P. Kawa and P. Syga received support from the Department of Artificial Intelligence, Wrocław University of Science and Technology. P. Spurek and M. Rut were supported by the National Center of Science (Poland) Grant No. 2023/50/E/ST6/00068.
@@ A.1. Extended ablation study
Length $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
Compared to the 1.5 s case, where a spline order of 3 is clearly optimal, its advantage is less evident in Tab. 14. KAN with a spline order of 5 attains higher PSNR and SI-SNR, while remaining competitive in PESQ and STOI.
@@
Order $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
The grid size results (cf. Tab. 15) show trends similar to the 1.5 s case (Tab. 10), increasing grid size improves all metrics except LSD.
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
Rate $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
Spline $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$